1. An electrically neutral glass rod is rubbed against a neutral piece of silk. During this process 2.0 x 1015 electrons are transferred from the rod to the silk.
a. How much charge (in coulombs) did the silk acquire?

b. What charge does the glass rod have after being rubbed on the silk?
1
Expert's answer
2020-04-27T10:00:46-0400

The elemental electric charge corresponds to "e=1.60\cdot 10^{-19}C". The electron charge of the electron is assigned a negative sign. The electric charge of "N=2.0\cdot 10^{15}" electrons corresponds to "Q=-e\cdot N=-3.20\cdot 10^{-4}C". In nature, the law of conservation of electric charge applies, therefore if electrons pass from neutral glass rod to the silk then glass rod turns out to be positively charged, and silk to be negatively charged.

Answer: a. The silk acquire charge "Q=-3.20\cdot 10^{-4}C" .

b.  The glass rod have an charge "Q=3.20\cdot 10^{-4}C" after being rubbed on the silk.